### v1

@@ Abstract
Atypical mitoses mark a deviation in the cell division process that can be an independent prognostically relevant marker for tumor malignancy. However, their identification remains challenging due to low prevalence, at times subtle morphological differences from normal mitoses, low inter-rater agreement among pathologists, and class imbalance in datasets. Building on the Atypical Mitosis dataset for Breast Cancer (AMi-Br), this study presents a comprehensive benchmark comparing deep learning approaches for automated atypical mitotic figure (AMF) classification, including baseline models, foundation models with linear probing, and foundation models fine-tuned with low-rank adaptation (LoRA). For rigorous evaluation, we further introduce two new hold-out AMF datasets - AtNorM-Br, a dataset of mitoses from the The TCGA breast cancer cohort, and AtNorM-MD, a multi-domain dataset of mitoses from the MIDOG++ training set. We found average balanced accuracy values of up to 0.8135, 0.7696, and 0.7705 on the in-domain AMi-Br and the out-of-domain AtNorm-Br and AtNorM-MD datasets, respectively, with the results being particularly good for LoRA-based adaptation of the Virchow-line of foundation models. Our work shows that atypical mitosis classification, while being a challenging problem, can be effectively addressed through the use of recent advances in transfer learning and model fine-tuning techniques. We make available all code and data used in this paper in this github repository: https://github.com/DeepMicroscopy/AMi-Br_Benchmark.
@@
Atypical Mitoses, Deep Learning, Foundation Models, Classification, Benchmarking, Histopathology, low-rank adaptation
@@ 1 Introduction
Mitosis is the process where a cell replicates its genetic material (DNA) and then divides into two identical daughter cells. This is how cells grow, repair damaged tissues, and replace old or worn-out cells. The cell in the process of dividing to create two new cells may be observed microscopically as a mitotic figure (MF). Quantifying tumor proliferation is essential for assessing many human and animal cancers, including human breast cancer (Fitzgibbons and Connolly, 2023; Kiupel et al., 2011; Louis et al., 2016). A key metric in this assessment is the mitotic count, which reflects the number of cells in mitosis and provides valuable information about the tumor’s growth rate and, consequently, its histological grade. Elevated mitotic counts are associated with an increased risk of distant metastasis and reduced patient survival (Medri et al., 2003). Identification of MFs largely remains a manual task, making it highly subjective and time intensive. This task is furthermore subject to a high inter-rater variability, complicating the creation of datasets used for automatic detection (Meyer et al., 2005, 2009; Wilm et al., 2021). Deep learning models have been extensively used in tasks like mitosis detection and classification (Aubreville et al., 2024; Veta et al., 2019; Aubreville et al., 2023b), and they offer a promising outlook on the problem.
@@
In tumor samples, we find two variations of MFs: normal and atypical. Normal MFs follow the typical, orderly process of cell division, including stages like prometaphase, metaphase, anaphase, and telophase, where chromosomes align and separate to produce two identical daughter cells Donovan et al. (2021). An atypical mitotic figure (AMF), on the other hand, is a cell undergoing an abnormal cell division process characterized by errors in chromosome segregation or other morphological irregularities, often seen in cancer cells (Donovan et al., 2021). The rate of AMFs has been correlated with poorer prognosis for breast cancer (Ohashi et al., 2018; Lashen et al., 2022) and other tumor types (Jin et al., 2007; Kalatova et al., 2015; Bertram et al., 2023; Matsuda et al., 2016). Recently, Jahanifar et al. (2025) have confirmed the prognostic relevance of the rate and density of AMFs across a wide range of tumors in a large-scale evaluation.
Yet, the identification of AMF is complicated by a low inter-rater agreement for classifying the mitoses into normal or atypical classes, due to challenging and irregular morphologies (Bertram et al., 2023; Aubreville et al., 2023a). Deep learning models have the potential to uniformly detect and classify mitoses, enable reproducibility, and reduce the workload of pathologists in this regard. Additionally, automatic, deep learning-based analysis allows for the large-scale, data-driven identification of prognostically relevant markers, such as the ratio of AMFs on the whole slide (Jahanifar et al., 2025).
@@
There have been very few studies investigating deep learning for AMFs (Aubreville et al., 2023a; Fick et al., 2024; Bertram et al., 2025) The classification of atypical vs. normal MFs is complicated by the overlapping morphologies of the normal and atypical mitotic figures (Donovan et al., 2021), and the low frequency of AMF, resulting in class imbalance. Although this is a prognostically relevant problem, there is currently a lack of standardized benchmarks, which makes it difficult to compare approaches, hence it is crucial to develop suitable datasets and establish strong baselines that reflect the unique challenges of this task. Recently, foundation models, which are mostly vision transformer (ViT)-based models trained on millions of histopathology slides using self supervised learning, have gained popularity in medical imaging tasks like classification. These models are typically employed either via linear probing, where a lightweight classifier is trained on top of a frozen backbone, or through parameter-efficient fine-tuning techniques like Low Rank Adaptation (LoRA) (Hu et al., 2022). But there is a lack of comprehensive studies on using these methods for atypical vs. normal mitoses classification.
@@
The contribution of this work is two-fold: First, we present three datasets, centered around the identification of AMFs, and secondly, we provide a comprehensive benchmark comparing widely used deep learning architectures, foundation models with linear probing, and foundation models fine-tuned with LoRA for the task of atypical versus normal mitosis classification.
@@ 2 Datasets
We present and use three distinct datasets in this paper - AMi-Br, AtNorM-Br and AtNorM-MD, as described below. This work is an extension of a previous conference contribution by our group (Bertram et al., 2025), which already introduced the first of those datasets (AMi-Br) in brief.
Being the largest scale dataset of the set, we use data from the AMI-Br dataset for training in all evaluations. This allows us to use the other two – newly introduced – datasets as independent hold out sets. It has been shown that data distribution shifts (domain shift) impede the performance of deep learning models significantly (Stacke et al., 2020; Aubreville et al., 2021). With these datasets, we can expect varying degrees of domain shifts, as they stem from different sources (AtNorm-Br) or even tumor types (AtNorm-MD), which allows us to investigate generalization. The statistics of the dataset, including total number of atypical and normal mitotic figures in each, annotation type, source dataset(s), species involved etc. are summarized in Table 1.
@@ 2.1 AMi-Br
The first dataset, AMi-Br (Bertram et al., 2025) consists of a total of 3,720 MFs from human breast cancer, of which 1,999 MFs are from the TUPAC16 (Veta et al., 2019) alternative label set (Bertram et al., 2020) and 1,721 MFs are from the MIDOG21 training dataset (Aubreville et al., 2023b). The MFs were classified into one of two classes — atypical or normal — by three expert pathologists. According to the result of the majority vote between the three experts, 832 MFs were found to be atypical, while the remaining 2,888 were classified as normal MFs. We found total agreement per object by all three experts in 2908 (78.17%) of the cases.
@@ 2.2 AtNorM-MD
The second dataset, Atypical and Normal Mitosis (AtNorM)-MultiDomain (MD) extends beyond human breast cancer and covers six domains, spanning both human and canine tumors. These include canine lung cancer, canine lymphoma, canine cutaneous mast cell tumor, human neuroendocrine tumor, canine soft tissue sarcoma, and human melanoma. It was sourced from all but the human breast cancer domains of the publicly available MIDOG++ dataset (Aubreville et al., 2023c). It comprises 2,107 mitotic figures, with 219 (10.4%) being atypical. Labeling was performed as a majority vote of five pathology experts, three of which were board-certified. We found total agreement per object by all five experts in 1,466 (69.58%) of the cases.
@@ 2.3 AtNorM-Br
The third dataset, AtNorM-Br, also made available publicly within the scope of this work, contains 746 MF instances from 179 patients from the breast cancer (BRCA) cohort of the The Cancer Genome Atlas (TCGA) (Lingle et al., 2016). TCGA contains images from various sources and with partially mixed quality and is thus also a valuable asset for the assessment of generalization. This dataset was annotated by a single expert with high experience in AMF classification, according to which 128 MFs were found to be atypical and the remaining 618 were classified as normal, yielding an AMF rate of 17.16%.
@@
Model Pretraining Algorithm Training Dataset Size Model Type Size (Params) WSIs Tiles UNI DINOv2 100K 100M ViT-L/16 307M UNI2-h DINOv2 350K 200M ViT-H/14 681M Virchow DINOv2 1.5M 2B ViT-H/14 632M Virchow2 DINOv2 3.4M 1.7B ViT-H/14 632M Prov-Gigapath DINOv2 170K 1.3B ViT-g/16 1.1B H-Optimus-0 DINOv2 500K - ViT-g/14 1.1B H-Optimus-1 DINOv2 1M 2B ViT-g/14 1.13B
@@ 3.1 End-to-end-trained Deep Learning Models
To establish robust baselines for atypical vs. normal mitoses classification, we evaluate three widely-used deep learning architectures – EfficientNetV2 (Tan and Le, 2021), ViT (Dosovitskiy et al., 2020), and Swin Transformer (Liu et al., 2021). All of these models have demonstrated strong performance across various medical imaging datasets and tasks. EfficientNetV2 is a well-established convolutional neural network (CNN) known for its strong performance and efficiency in various classification tasks. It serves as a representative of traditional CNN-based approaches. In contrast, the ViT represents a paradigm shift from convolutional architectures to transformer-based models. A ViT processes images as sequences of patches and leverages self-attention mechanisms to model global relationships across the input. This global receptive field allows the ViT to capture long-range dependencies, which can be particularly useful in complex histopathological patterns. The Swin Transformer further advances the transformer-based approach by introducing a hierarchical architecture with shifted windows. This design enables the model to compute attention locally while gradually building up to global representations, providing a balance between computational efficiency and the ability to capture both local details and global structure. Swin Transformers have been particularly effective in dense prediction tasks and high-resolution image analysis, making it well-suited for mitotic figure classification (Liu et al., 2021). Moreover, the selected model architectures were also chosen to be of similar size. Each model was trained in an end-to-end manner and evaluated on three held-out test datasets: a random, case-based split of the AMi-Br dataset, accounting for $\approx$ 22 % of the samples in the dataset (samples not used in training or validation), and the entire AtNorM-MD and AtNorM-Br datasets.
@@ 3.2 Foundation Models
Foundation models are models that are trained on large and diverse corpora of data, often using unsupervised techniques like self supervision. The goal of training is to train feature extractors that generalize well and are easily adaptable to down-stream tasks. In the field of computational pathology, we have seen a very strong incline in publicly available models, warranting an analysis on the task of AMF classification. We compare eight state-of-the-art models – UNI (Chen et al., 2024), UNI2-h (Chen et al., 2024; MahmoodLab, 2025), Virchow (Vorontsov et al., 2024), Virchow2 (Zimmermann et al., 2024), Prov-Gigapath (Xu et al., 2024), H-Optimus-0 (Saillard et al., 2024), H-Optimus-1 (Bioptimus, 2025), and H0-mini (Filiot et al., 2025). An overview of the details for each model is provided in Table 2. All of the selected foundation models are ViT-based and have been pre-trained on datasets comprising of hundreds of thousands to multiple millions of whole slide images.
@@ 3.2.1 Linear Probing of Foundation Models
We use a linear probing strategy, where the feature extractor (i.e., the foundation model) is used to extract the embeddings. After that, the foundation model is kept frozen, and a linear classifier is trained on top of the extracted features. This process helps us to assess the ability of the representations of the foundation models to classify atypical vs. normal mitoses.
@@ 4 Experimental Setup
We set up detailed experiments for the above cases. The image preprocessing for the binary classification experiments included resizing all patches to 224×224 pixels. Training augmentations consisted of random horizontal flip, rotation, color jitter, and random resized crop, followed by normalization with ImageNet statistics. Validation images were only resized and normalized. The best checkpoint for each fold was selected based on validation balanced accuracy. All experiments were performed using 5-fold cross validation, where stratification of the data was based on the slide (patient) to ensure that all patches from a given slide appeared in only one split (train or validation), thus preventing data leakage. To mitigate class imbalance, we employed weighted random sampling during training, with sample weights inversely proportional to class frequencies. All models were trained using the Adam optimizer with L2 regularization (learning rate: 1e-4, weight decay: 1e-5) and cross-entropy loss. We applied early stopping with a patience of 15 epochs based on validation balanced accuracy, and used learning rate scheduling on plateau (factor: 0.5, patience: 3, minimum LR: 1e-7). Each model was trained for up to 100 epochs with a batch size of 8. The rest of the case-specific details are described below in the subsections.

### v2

@@ Abstract
Atypical mitosis marks a deviation in the cell division process that has been shown be an independent prognostic marker for tumor malignancy. However, atypical mitosis classification remains challenging due to low prevalence, at times subtle morphological differences from normal mitotic figures, low inter-rater agreement among pathologists, and class imbalance in datasets. Building on the Atypical Mitosis dataset for Breast Cancer (AMi-Br), this study presents a comprehensive benchmark comparing deep learning approaches for automated atypical mitotic figure (AMF) classification, including end-to-end trained deep learning models, foundation models with linear probing, and foundation models fine-tuned with low-rank adaptation (LoRA). For rigorous evaluation, we further introduce two new held-out AMF datasets - AtNorM-Br, a dataset of mitotic figures from the TCGA breast cancer cohort, and AtNorM-MD, a multi-domain dataset of mitotic figures from a subset of the MIDOG++ training set. We found average balanced accuracy values of up to 0.8135, 0.7788, and 0.7723 on the in-domain AMi-Br and the out-of-domain AtNorm-Br and AtNorM-MD datasets, respectively. Our work shows that atypical mitotic figure classification, while being a challenging problem, can be effectively addressed through the use of recent advances in transfer learning and model fine-tuning techniques. We make all code and data used in this paper available in this github repository: https://github.com/DeepMicroscopy/AMi-Br_Benchmark.
@@
Atypical Mitosis, Deep Learning, Foundation Models, Classification, Benchmarking, Histopathology, low-rank adaptation
@@ 1 Introduction
Mitosis is the process where a cell replicates its genetic material (DNA) and then divides into two identical daughter cells. This process enables the proliferation of cells, supporting physiological tissue growth and facilitating the replacement of old, damaged, or worn-out cells. The dividing cell may be observed microscopically as an mitotic figure (MF) with specific morphologies resembling the highly regulated cell division phases. However, cell proliferation does not occur only under physiological conditions; it is also a hallmark of cancer, where cell division is dysregulated and increased. Subsequently, quantifying cell proliferation is essential for estimating the aggressiveness of many human and animal cancers, including human breast cancer (Fitzgibbons and Connolly, 2023; Kiupel et al., 2011; Louis et al., 2016). The routine method in the assessment of the growth rate is the mitotic count, which reflects the number of MFs within a specific tumor area. Elevated mitotic counts are associated with an increased risk of distant metastasis and higher risk of death (Medri et al., 2003).
@@
In tumor samples, two variants of MFs can be found: normal and atypical. Normal MFs indicate the orderly process of cell division, designated by the prometaphase, metaphase, anaphase, and telophase, where chromosomes align and separate equally to produce two identical daughter cells (Donovan et al., 2021). An atypical mitotic figure (AMF), on the other hand, is a cell undergoing an abnormal cell division process characterized by errors in chromosome segregation or other morphological irregularities, which promote cancer progression (Donovan et al., 2021). Until recently, a correlation of the count of AMFs with shorter patient survival has been demonstrated in only few studies on breast cancer (Ohashi et al., 2018; Lashen et al., 2022) and selected other tumor types (Jin et al., 2007; Kalatova et al., 2015; Bertram et al., 2023; Matsuda et al., 2016). A very recent study by Jahanifar et al. (2025) has expanded these insights and the corresponding prognostic relevance of the rate of AMFs across a wide range of tumors in a large-scale evaluation, highlighting the importance of further investigations of this prognostic parameter.
Identification of MFs is traditionally done by pathologists, making it highly subjective and time intensive. This process is further complicated by inter-observer variability, low prevalence of MFs, and inconsistencies in slide quality (Meyer et al., 2005, 2009; Wilm et al., 2021). To address these challenges, deep learning-based MF detection has been extensively investigated (Aubreville et al., 2024; Veta et al., 2019; Aubreville et al., 2023b), demonstrating a promising outlook for improving both accuracy and efficiency and reducing the diagnostic workload for pathologists. Along similar lines, deep learning models offer potential for the automatic differentiation of AMFs against normal MFs. For instance, this could enable the quantification of the ratio of AMFs across entire whole slide images (Jahanifar et al., 2025), which is otherwise impractical to perform manually due to time constraints and the rarity of AMFs. Such quantitative assessments can support more objective and reproducible evaluations of tumor aggressiveness, potentially improving treatment planning in clinical settings.
@@
At the time of this writing, there have been very few studies investigating deep learning for AMFs (Aubreville et al., 2023a; Fick et al., 2024; Bertram et al., 2025). The automated classification of atypical vs. normal MFs is complicated by the substantial variation of morphologies, often with some mitotic figures displaying overlapping features of both normal and AMFs (Donovan et al., 2021; Bertram et al., 2023), and the low frequency of AMF, resulting in class imbalance. Given the prognostic relevance of differentiating normal vs. atypical MFs, standardized benchmarks, such as publicly available datasets, evaluation protocols, and baseline models, may serve as facilitators to push further developments in this field that tackle these challenges. In this context, foundation models offer a promising direction. These models, typically vision transformer (ViT)-based and trained on millions of histopathology slides using self supervised learning, have shown strong performance across various medical imaging tasks, including classification (Campanella et al., 2025; Breen et al., 2025), especially in small data regimes. These models are typically employed either via linear probing, where a lightweight classifier is trained on top of a frozen backbone, or through parameter-efficient fine-tuning techniques like Low Rank Adaptation (LoRA) (Hu et al., 2022). However, comprehensive studies leveraging these techniques for atypical vs. normal mitosis classification are still lacking.
@@
The contribution of this work is two-fold: First, we present three datasets, designed for the automated identification of AMFs to establish suitable datasets that reflect the unique challenges of this task. Second, we provide a comprehensive benchmark comparing end-to-end trained deep learning architectures, foundation models with linear probing, and foundation models fine-tuned with LoRA for the task of atypical versus normal mitotic figure classification.
@@ 2 Datasets
We present and use three distinct datasets in this paper - AMi-Br, AtNorM-Br, and AtNorM-MD as described below. This work is an extension of a previous conference contribution by our group (Bertram et al., 2025), which already introduced the first of those datasets (AMi-Br) in brief. Representative image samples of atypical and normal mitotic figures from all three datasets are shown in Figure 1.
Being the largest of the three datasets used in this study, we use data from the AMi-Br dataset for training in all evaluations. This allows us to use the other two newly introduced datasets as independent held-out sets. It has been shown that data distribution shifts (domain shifts) impede the performance of deep learning models significantly (Stacke et al., 2020; Aubreville et al., 2021). With these datasets, we can expect varying degrees of domain shifts, and which allows us to investigate generalization performance. Only a part of the AMi-Br dataset is used for training the models for binary classification of atypical vs. normal mitoses, and the rest is used as another held-out test dataset. The dataset statistics, including total number of atypical and normal MFs in each, annotation type, source dataset(s), species involved etc. have been summarized in Table 1.
@@ 2.1 AMi-Br
The first dataset, AMi-Br (Bertram et al., 2025) consists of a total of 3,720 MFs from human breast cancer, of which 1,999 MFs are from the TUPAC16 (Veta et al., 2019) alternative label set (Bertram et al., 2020) and 1,721 MFs are from the MIDOG21 training dataset (Aubreville et al., 2023b). The MFs were graded into one of two classes — atypical or normal — by three expert pathologists, two of whom were board-certified. According to the result of the majority vote between the three experts, 832 MFs were found to be atypical, while the remaining 2,888 were classified as normal MFs. We found total agreement per object by all three experts in 2908 (78.2%) of the cases.
@@ 2.2 AtNorM-MD
The second dataset, Atypical and Normal Mitosis (AtNorM)-MultiDomain (MD) extends beyond human breast cancer and covers six domains, spanning both human and canine tumors. These include canine lung cancer, canine lymphoma, canine cutaneous mast cell tumor, human neuroendocrine tumor, canine soft tissue sarcoma, and human melanoma. It was sampled randomly from all but the human breast cancer domains of the publicly available MIDOG++ dataset (Aubreville et al., 2023c). It comprises 2,107 MFs from 70 patients, with 219 (10.4%) being atypical. Labeling was performed as a majority vote of five pathology experts, three of whom were board-certified. We found total agreement per object by all five experts in 1,466 (69.6%) of the cases.
@@ 2.3 AtNorM-Br
The third dataset, AtNorM-Br, also made available publicly within the scope of this work, contains 746 MF instances from 179 patients from the breast cancer (BRCA) cohort of the The Cancer Genome Atlas (TCGA) (Lingle et al., 2016). TCGA contains images from various sources and with partially mixed quality, stemming from differences in staining protocols, scanning equipment, and tissue preparation across institutions, and is thus also a valuable asset for the assessment of generalization. This dataset was annotated by a single expert with high experience in AMF classification, according to which 128 MFs were found to be atypical and the remaining 618 were classified as normal, yielding an AMF rate of 17.16%.
@@
Model Training Dataset Size Model Type Size (Params) WSIs Tiles UNI 100K 100M ViT-L/16 307M UNI2-h 350K 200M ViT-H/14 681M Virchow 1.5M 2B ViT-H/14 632M Virchow2 3.4M 1.7B ViT-H/14 632M Prov-Gigapath 170K 1.3B ViT-g/16 1.1B H-Optimus-0 500K - ViT-g/14 1.1B H-Optimus-1 1M 2B ViT-g/14 1.13B
@@ 3.1 End-to-end-trained Deep Learning Models
To establish robust baselines for atypical vs. normal mitoses classification, we evaluate three widely-used deep learning architectures – EfficientNetV2 (Tan and Le, 2021), ViT (Dosovitskiy et al., 2020), and Swin Transformer (Liu et al., 2021). All of these models have demonstrated strong performance across various medical imaging datasets and tasks. EfficientNetV2 is a well-established convolutional neural network (CNN) known for its strong performance and efficiency in various classification tasks. It serves as a representative of traditional CNN-based approaches. In contrast, the ViT represents a major shift from convolutional architectures to transformer-based models. A ViT processes images as sequences of patches and uses self-attention mechanisms to model global relationships across the input. This global receptive field allows the ViT to capture long-range dependencies, which can be particularly useful for complex histopathological patterns. The Swin Transformer further advances the transformer-based approach by introducing a hierarchical architecture with shifted windows. This architecture enables the model to compute attention locally while gradually building up to global representations, providing a balance between computational efficiency and the ability to capture both local details and global structure. Swin Transformers have been particularly effective in dense prediction tasks and high-resolution image analysis, making it well-suited for mitotic figure classification (Liu et al., 2021). The selected model architectures were chosen to be of similar size in parameters. Each model was trained in an end-to-end manner and evaluated on three held-out test datasets: a random, patient-stratified split of the AMi-Br dataset, accounting for $\approx$ 22 % of the samples in the dataset (samples not used in training or validation), and the entire AtNorM-MD and AtNorM-Br datasets.
@@ 3.2 Foundation Models
Foundation models are models that are trained on large and diverse corpora of data, often using unsupervised techniques like self supervision. The goal of foundation models is to train feature extractors that generalize well and are easily adaptable to down-stream tasks. In the field of computational pathology, we have seen a very strong incline towards publicly available foundation models, calling for an investigation on the task of AMF classification. We compare eight state-of-the-art models – UNI (Chen et al., 2024), UNI2-h (Chen et al., 2024; MahmoodLab, 2025), Virchow (Vorontsov et al., 2024), Virchow2 (Zimmermann et al., 2024), Prov-Gigapath (Xu et al., 2024), H-Optimus-0 (Saillard et al., 2024), H-Optimus-1 (Bioptimus, 2025), and H0-mini (Filiot et al., 2025). An overview of selected details for each model such as training dataset size, size in parameters and model type for training is provided in Table 2. All selected foundation models are ViT-based and have been pre-trained using the DINOv2 algorithm (Oquab et al., 2023) on datasets comprising of hundreds of thousands to multiple millions of whole slide images.
@@ 3.2.1 Linear Probing of Foundation Models
We use a linear probing strategy, where the feature extractor (i.e., the foundation model) is used to extract the embeddings. After that, the foundation model is kept frozen, and a linear classifier is trained on top of the extracted features. This process helps us to assess the ability of the representations of the foundation models to classify atypical vs. normal MFs.
@@ 4 Experimental Setup
We set up detailed experiments for the above cases. The image preprocessing for the binary classification experiments included resizing all patches to 224×224 pixels. Training augmentations consisted of random horizontal flip, rotation, color jitter, and random resized crop, followed by normalization with ImageNet statistics. Validation images were only resized and normalized. The best checkpoint for each fold was selected based on validation balanced accuracy, which was used as the primary metric because of the class imbalance between the atypical and normal MFs in the dataset. All experiments were performed using 5-fold cross validation, where stratification of the data was based on the slide (patient) to ensure that all patches from a given slide appeared in only one split (train or validation), thus preventing data leakage. To mitigate class imbalance, we employed weighted random sampling during training, with sample weights inversely proportional to class frequencies. All models were trained using the Adam optimizer with L2 regularization (learning rate: $5\times 10^{-5}$, weight decay: $1\times 10^{-5}$) and cross-entropy loss. We applied early stopping with a patience of 15 epochs based on validation balanced accuracy, and used learning rate scheduling on plateau (factor: 0.5, patience: 3, minimum LR: $1\times 10^{-7}$). Each model was trained for up to 100 epochs with a batch size of 8. The rest of the case-specific details are described below in the subsections.